OpenStudy (anonymous):

The average of 6,7,5,8,2 and k is 8. What is the median of this set? A)8 B)6.5 C)6.75 D)5

OpenStudy (anonymous):

as mean is equal to (6+7+5+8+2+ k)/6 =8 (28 +k)/6 = 8 28+k = 48 k = 20 now you got the set of data 6,7,5,8,2,20 now apply median formula to find the median

OpenStudy (anonymous):

put the numbers in numerical order and find the middle number. If there is two middle numbers, add those two numbers together and divide by 2 to find the median.

OpenStudy (anonymous):

whoohoo!

OpenStudy (anonymous):

Thankyou rizwan and kellie :)

OpenStudy (anonymous):

thats what we are here for :)

OpenStudy (anonymous):

:)